OpenStudy (anonymous):

Solve the logarithmic equation algebraically. Approximate the result to three decimal places. (If there is no solution, enter NO SOLUTION.) log2X+log2(x-3)=2

OpenStudy (anonymous):

\[\log_{2}x+\log_{2}(x-3)=2 \]

OpenStudy (anonymous):

sorry about weird looking format on initial question.. The above is the correct format

jhonyy9 (jhonyy9):

so do you know the property of logarithms ?

jhonyy9 (jhonyy9):

how you can simplifie it log A +log B = ?

jhonyy9 (jhonyy9):

using property of logarithms

jhonyy9 (jhonyy9):

so than you need to learn it again than log a + log b = log a*b ok ?

jhonyy9 (jhonyy9):

suppose being base of 2

jhonyy9 (jhonyy9):

so and using log base 2 how you can writing the number 2 ?

OpenStudy (anonymous):

would it be \[2\ln +2\ln (x-3)=2\] to start ?

jhonyy9 (jhonyy9):

not is necessary because there is speciffied that us logarithm base 2

OpenStudy (anonymous):

\[2\log_{2}x+2\log_{2}(x-3)=2 \] is that correct to start process?

jhonyy9 (jhonyy9):

not is so there is log2 x + log2 (x-3) =2 yes ?

jhonyy9 (jhonyy9):

using logarithm base 2 you can writing the 2 how ?

jhonyy9 (jhonyy9):

so like 2= log2 2^2 right ?

jhonyy9 (jhonyy9):

than because you know that log2 a + log2 b = log2 a*b so using this property your exercise will be log2 x +log2 (x-3) =log2 2^2 log2 (x*(x-3)) = log2 2^2 so than result x*(x-3)=2^2 x^2 -3x =4 x^2 -3x -4=0 factorized will get (x-4)(x+1)=0 so than x-4=0 x=4 and x-1=0 x=1 so but you know that the function of logarithm is defined just in interval of (0,+infinite) so from this result that the value x=1 not is right because there is log2 (x-3) where using x=1 you will get log2 of -2 what is undefined so than will be right just x=4 hope so much that i have wrote it right,correct and understandably easy right good luck bye